# Grid-Agent: An LLM-Powered Multi-Agent System for Power Grid Control

### v2

@@ V-D Experimental Results and Analysis
Figure 2: Overall performance comparison across six LLMs. (a) Success rate shows the reliability of each model. (b) Average runtime indicates computational efficiency. (c, d) Average iterations and actions measure the conciseness and quality of the generated solutions.
As shown in Fig.2a, the more capable and efficient models, gemini-2.5-flash and gpt-4.1-mini, achieved a perfect 100% success rate across all scenarios. The flagship models, gemini-2.5-pro and gpt-4.1, also performed robustly with an 100% success rate. The smaller models, gemini-2.5-flash-lite and gpt-4.1-nano, struggled with the more complex scenarios, achieving 63% and 44% success rates, respectively.
There is a clear trade-off between performance and speed. Fig.2b shows that gemini-2.5-flash was by far the fastest model, resolving violations in under 6 seconds on average, whereas the larger gpt-4.1 and gemini-2.5-pro models required over 16 seconds. Figs.2c and 2d show that gemini-2.5-flash also found the most efficient solutions, requiring the fewest iterations and actions, indicating a strong inherent capability for structured, multi-step tool-use reasoning.
Figure 3: Comparison of action efficiency, measured as the average number of violations resolved per control action for successful runs. Higher values indicate more effective, coordinated solutions.
@@
Figure 4: Analysis of action strategies. (a) Heatmap showing the proportional usage of each action type. (b) Boxplot of coordination scores, indicating the quality of multi-action plans. (c) Violin plot showing the distribution of total actions required per solution. (d) Scatter plot correlating coordination score with success rate.
@@
Fig.4b shows the distribution of coordination scores, a metric assessing how well actions within a single step are logically grouped to address related violations. Models like gemini-2.5-pro and gemini-2.5-flash-lite consistently achieved high scores, indicating an ability to formulate synergistic action plans. The violin plot in Fig.4c further explores solution efficiency, showing that most successful models converged in under 10 actions. The wide distribution for gemini-2.5-flash-lite suggests it sometimes resorted to less efficient, sequential corrections.
Crucially, Fig. 4d demonstrates a strong positive correlation between the average coordination score and the final success rate. The top-performing models, gemini-2.5-flash and gpt-4.1-mini, are clustered in the top-right quadrant, confirming that the ability to generate well-coordinated, strategic plans is a key determinant of success in this complex problem domain. This validates that the LLM is not merely guessing but is performing effective, structured reasoning.
@@
Figure 5: Performance analysis across different network types. (a) Success rate heatmap by network category. (b) Average runtime heatmap by network category. (c) Boxplot of action efficiency, highlighting high performance on larger radial systems. (d) Scatter plot of runtime versus network size.
@@ VI Discussion
The experimental results demonstrate that Grid-Agent is a potent framework for resolving complex power grid violations. This section interprets these findings, discusses their implications for power systems operations, and considers the critical aspects of scalability, safety, and reliability.
VI-A Implications for Power Systems Operations
Grid-Agent demonstrates the potential for LLM-based systems to enhance power grid operations through intelligent automation combined with human-interpretable reasoning. The system’s ability to provide clear, step-by-step explanations for its decisions addresses a critical need in power systems operations, where understanding the rationale behind control actions is essential for operator confidence and regulatory compliance.
@@ VI-A Implications for Power Systems Operations
The coordinated approach to violation resolution, as evidenced by the high action efficiency (Fig.3), offers significant advantages over traditional methods. In scenarios involving renewable energy integration, violations are often interconnected across the network. Grid-Agent’s capacity to identify and execute strategic, high-impact actions—such as a single switch operation that resolves multiple downstream issues—is a marked improvement over myopic algorithms that address violations sequentially. This holistic reasoning is crucial for maintaining stability in increasingly dynamic and complex grids.
VI-B Scalability Considerations
The adaptive multi-scale representation framework directly addresses one of the primary challenges in applying LLMs to large-scale power systems: the limited context window. By dynamically adjusting the level of detail provided to the LLM based on network size and complexity, the system maintains practical computational requirements while preserving solution quality. The strong performance on the IEEE 69-bus system, particularly the superior efficiency of the gemini-2.5-flash model, validates this approach.
@@ VI-B Scalability Considerations
Future work should examine scalability to transmission-level networks with thousands of buses. Such large-scale applications will likely exceed the capabilities of a single-agent planner and may require further advances in hierarchical network representation and distributed multi-agent coordination, where regional agents manage local violations and communicate high-level state changes to a central coordinating agent. The integration of retrieval-augmented generation (RAG) will also be explored. Instead of serializing large network sections into the prompt, the LLM would query a specialized vector database containing detailed network topology, component specifications, and a history of successful action sequences for similar faults. This approach would significantly reduce the context size passed to the LLM, leading to faster inference, lower operational costs, and the ability to reason over vast historical and real-time data that would otherwise be computationally prohibitive.
Furthermore, while this exploratory study has focused on the feasibility and strategic reasoning capabilities of Grid-Agent, a crucial direction for future work is the formal analysis of solution optimality. We plan to conduct rigorous benchmarking studies comparing the control actions proposed by Grid-Agent against traditional optimization techniques, particularly mixed-integer nonlinear Optimal Power Flow (OPF) solvers. Such a comparison will be essential to quantify the trade-offs between the LLM’s heuristic, rapid, and interpretable decision-making and the mathematically guaranteed optimality of conventional methods. This will help establish clear performance envelopes for where an agentic approach excels and where traditional optimizers remain indispensable, potentially leading to hybrid systems that leverage the strengths of both.
VI-C Safety and Reliability
The comprehensive validation and rollback mechanisms are cornerstones of the Grid-Agent framework, providing essential safety guarantees for critical infrastructure applications. The system’s ability to explore potential solutions within a secure sandbox, verify their efficacy, and automatically revert failed attempts ensures that the operational grid is never exposed to untested or detrimental actions. This demonstrates the feasibility of AI-driven optimization in safety-critical domains.
However, deployment in a live operational environment would require additional safeguards. These should include robust human-in-the-loop oversight, staged deployment protocols (e.g., operating in an advisory-only mode before granting control), and comprehensive testing under a wider diversity of fault conditions and adversarial scenarios to ensure resilience and build operational trust.
VI-D Integrating Reinforcement Learning for Time-Series Control
While Grid-Agent excels at formulating discrete, strategic actions, a key avenue for future improvement is the integration with Reinforcement Learning (RL) agents for optimizing continuous, time-series control problems. In this paradigm, the LLM-based Planner would act as a high-level orchestrator, identifying the need for a dynamic response (e.g., ”Utilize battery storage to smooth a predicted renewable generation drop over the next hour”).
This high-level directive would then be passed to a specialized RL agent trained to generate an optimal, time-dependent action schedule. This would enable Grid-Agent to support complex, continuous control tasks such as generating a precise State of Charge (SOC) trajectory for battery systems or a dynamic load curtailment schedule over a given time horizon. This hybrid approach would combine the LLM’s semantic reasoning and strategic planning with the RL agent’s proficiency in high-dimensional, sequential decision-making, moving beyond single-shot corrections to proactive, continuous grid management.
@@ VII Conclusion
This paper presented Grid-Agent, a novel multi-agent system that successfully integrates the semantic reasoning of Large Language Models with the numerical precision of power systems optimization. The system demonstrates significant improvements in violation resolution efficiency and provides human-interpretable explanations, which are essential for modern power systems operations.
Key achievements of this work include the development of a modular multi-agent architecture for decomposed problem-solving, an adaptive multi-scale representation framework that enables LLM application to larger networks, and a coordinated action optimization strategy that resolves multiple violations with minimal intervention. The integrated safety mechanisms, including sandboxed execution and automated rollbacks, ensure that all proposed solutions are validated before affecting the system, making the framework suitable for critical infrastructure applications.
The experimental validation on standard IEEE and CIGRE test networks confirms the system’s effectiveness across diverse violation scenarios, with particular strengths in complex, multi-violation situations that challenge traditional optimization approaches. The built-in training data generation framework provides a foundation for continuous learning and adaptation, enabling the system to improve over time. Grid-Agent represents a significant step toward building intelligent, adaptive, and trustworthy power grid management systems that combine the reasoning capabilities of large language models with the rigorous requirements of critical infrastructure.
Acknowledgment
The authors acknowledge the support of the open-source community, particularly the developers of pandapower [18] and LangGraph frameworks that enabled this research.

### v3

@@ V-D Experimental Results and Analysis
Figure 2: Overall performance comparison across six LLMs, (a) Success rate shows the reliability of each model, (b) Average runtime indicates computational efficiency. (c, d) Average iterations and actions measure the conciseness and quality of the generated solutions
As shown in Fig.2a, the more capable and efficient models, gemini-2.5-flash and gpt-4.1-mini, achieved a perfect 100% success rate across all scenarios. The flagship models, gemini-2.5-pro and gpt-4.1, also performed robustly with an 100% success rate. The smaller models, gemini-2.5-flash-lite and gpt-4.1-nano, struggled with the more complex scenarios, achieving 63% and 44% success rates, respectively. There is a clear trade-off between performance and speed. Fig.2b shows that gemini-2.5-flash was by far the fastest model, resolving violations in under 6 seconds on average, whereas the larger gpt-4.1 and gemini-2.5-pro models required over 16 seconds. Figs.2c and 2d show that gemini-2.5-flash also found the most efficient solutions, requiring the fewest iterations and actions, indicating a strong inherent capability for structured, multi-step tool-use reasoning.
Figure 3: Comparison of action efficiency, measured as the average number of violations resolved per control action for successful runs. Higher values indicate more effective, coordinated solutions
@@
Figure 4: Analysis of action strategies, (a) Heatmap showing the proportional usage of each action type, (b) Boxplot of coordination scores, indicating the quality of multi-action plans, (c) Violin plot showing the distribution of total actions required per solution, (d) Scatter plot correlating coordination score with success rate
@@
Fig.4b shows the distribution of coordination scores, a metric assessing how well actions within a single step are logically grouped to address related violations. Models like gemini-2.5-pro and gemini-2.5-flash-lite consistently achieved high scores, indicating an ability to formulate synergistic action plans. The violin plot in Fig.4c further explores solution efficiency, showing that most successful models converged in under 10 actions. The wide distribution for gemini-2.5-flash-lite suggests it sometimes resorted to less efficient, sequential corrections. Crucially, Fig. 4d demonstrates a strong positive correlation between the average coordination score and the final success rate. The top-performing models, gemini-2.5-flash and gpt-4.1-mini, are clustered in the top-right quadrant, confirming that the ability to generate well-coordinated, strategic plans is a key determinant of success in this complex problem domain. This validates that the LLM is not merely guessing but is performing effective, structured reasoning.
@@
Figure 5: Performance analysis across different network types, (a) Success rate heatmap by network category, (b) Average runtime heatmap by network category, (c) Boxplot of action efficiency, highlighting high performance on larger radial systems, (d) Scatter plot of runtime versus network size
@@ VI Discussion
Grid-Agent demonstrates the potential for LLM-based systems to enhance power grid operations through intelligent automation combined with human-interpretable reasoning. The system’s ability to provide clear, step-by-step explanations for its decisions addresses a critical need in power systems operations, where understanding the rationale behind control actions is essential for operator confidence and regulatory compliance. The coordinated approach to violation resolution, as evidenced by the high action efficiency (Fig.3), offers significant advantages over traditional methods. In scenarios involving renewable energy integration, violations are often interconnected across the network. Grid-Agent’s capacity to identify and execute strategic, high-impact actions—such as a single switch operation that resolves multiple downstream issues—is a marked improvement over myopic algorithms that address violations sequentially. This holistic reasoning is crucial for maintaining stability in increasingly dynamic and complex grids.
@@
The adaptive multi-scale representation framework directly addresses one of the primary challenges in applying LLMs to large-scale power systems: the limited context window. By dynamically adjusting the level of detail provided to the LLM based on network size and complexity, the system maintains practical computational requirements while preserving solution quality. The strong performance on the IEEE 69-bus system, particularly the superior efficiency of the gemini-2.5-flash model, validates this approach. Future work should examine scalability to transmission-level networks with thousands of buses. Such large-scale applications will likely exceed the capabilities of a single-agent planner and may require further advances in hierarchical network representation and distributed multi-agent coordination, where regional agents manage local violations and communicate high-level state changes to a central coordinating agent. The integration of retrieval-augmented generation (RAG) will also be explored. Instead of serializing large network sections into the prompt, the LLM would query a specialized vector database containing detailed network topology, component specifications, and a history of successful action sequences for similar faults. This approach would significantly reduce the context size passed to the LLM, leading to faster inference, lower operational costs, and the ability to reason over vast historical and real-time data that would otherwise be computationally prohibitive. Furthermore, while this exploratory study has focused on the feasibility and strategic reasoning capabilities of Grid-Agent, a crucial direction for future work is the formal analysis of solution optimality. We plan to conduct rigorous benchmarking studies comparing the control actions proposed by Grid-Agent against traditional optimization techniques, particularly mixed-integer nonlinear Optimal Power Flow (OPF) solvers. Such a comparison will be essential to quantify the trade-offs between the LLM’s heuristic, rapid, and interpretable decision-making and the mathematically guaranteed optimality of conventional methods. This will help establish clear performance envelopes for where an agentic approach excels and where traditional optimizers remain indispensable, potentially leading to hybrid systems that leverage the strengths of both.
@@ VII Conclusion
This paper presented Grid-Agent, a novel multi-agent system that successfully integrates the semantic reasoning of Large Language Models with the numerical precision of power systems optimization. The system demonstrates significant improvements in violation resolution efficiency and provides human-interpretable explanations, which are essential for modern power systems operations. Key achievements of this work include the development of a modular multi-agent architecture for decomposed problem-solving, an adaptive multi-scale representation framework that enables LLM application to larger networks, and a coordinated action optimization strategy that resolves multiple violations with minimal intervention. The integrated safety mechanisms, including sandboxed execution and automated rollbacks, ensure that all proposed solutions are validated before affecting the system, making the framework suitable for critical infrastructure applications. The experimental validation on standard IEEE and CIGRE test networks confirms the system’s effectiveness across diverse violation scenarios, with particular strengths in complex, multi-violation situations that challenge traditional optimization approaches. The built-in training data generation framework provides a foundation for continuous learning and adaptation, enabling the system to improve over time. Grid-Agent represents a significant step toward building intelligent, adaptive, and trustworthy power grid management systems that combine the reasoning capabilities of large language models with the rigorous requirements of critical infrastructure.